# Perturbation-Robust Predictive Modeling of Social Effects by Network Subspace Generalized Linear Models

### v4

@@ 3 Statistical Inference Properties
This section provides theoretical results for the estimation consistency and statistical inference of the proposed method. To this end, we need the following regularity conditions.
@@
Assumption 3 (Boundedness of $X$ and $W$).
@@
There exists constants $\delta,C>0$ such that when $\|\gamma-\gamma^{*}\|<\delta$ the oracle information matrix defined in (13) satisfies that
@@
$\bar{\tau}_{n}:=n^{-3/2}\frac{\|(\tilde{W}\tilde{W}^{T}-WW^{T})Z\|}{\min\left\{\left(1-\sigma_{r+1}\right)^{3},\sigma_{r+s}^{3}\right\}},$
@@
$\bar{\tau}_{n}=o(n^{-1/2}).$
Assumption 5 is our essential requirement for the network perturbation level that is tolerable. This assumption is also used in Le and Li (2022), which is shown to be reasonable for network perturbations in many cases. In Section 4, we will show that this is valid to handle many real-world applications.
Assumption 6 (Moment Constraint for Responses).
There exists constants $c>0$, $M_{0}>0$ and $\xi>2$ such that
@@
Assumption 6 provides a sufficient condition for the Lindeberge-Feller Central Limit Theorem to hold. A similar constraint has been adopted in Yin et al. (2006); Gao et al. (2012).
@@
Consider the estimating equation (16) and assume that Assumptions 16 hold. For each $n$, there exists $\hat{\gamma}$ such that as $n\to\infty$,
@@
Theorem 1 shows that for each $n$, there exists a solution to the estimating equation (16) with high probability. In addition, the sequences of corresponding estimates for the true parameters in (3) are consistent. It is worth noting that similar to Yin et al. (2006), Theorem 1 itself does not guarantee the uniqueness of the solution $\hat{\gamma}$. This is because the log-likelihood function is generally not concave for special link functions. However, Corollary 2 below shows that restricting the model space to the class with natural links functions, or more generally, link functions ensuring concavity, leads to the uniqueness.
@@
Assume that Assumptions 1 to 6 hold and the link function is natural. That is, $h^{-1}=\psi$. Then the estimates in Theorem 1 are unique for sufficiently large $n$.
Our next result concerns the asymptotic distributions of the proposed estimates for $\theta^{*}$, $\beta^{*}$, and $\alpha^{*}$. Since these parameters depend on $\gamma^{*}$ through equations (9), (10), and (11), we need the covariance matrices of $\hat{\gamma}_{1:r}$, $\hat{\gamma}_{(r+1):p}$, and $\hat{\gamma}_{(p+1):(p+K-r)}$. These matrices can be estimated by the diagonal blocks of the inverse of the sample information matrix in (17), which we denote by $\tilde{F}^{-1}_{1}(\hat{\gamma})$, $\tilde{F}^{-1}_{2}(\hat{\gamma})$, and $\tilde{F}^{-1}_{3}(\hat{\gamma})$, respectively. Thus,
$\tilde{F}^{-1}(\hat{\gamma})$ $=\left(\begin{array}[]{ccc}\tilde{F}^{-1}_{1}(\hat{\gamma})&*&*\\ &\tilde{F}^{-1}_{2}(\hat{\gamma})&*\\ &*&\tilde{F}^{-1}_{3}(\hat{\gamma})\end{array}\right),$
where $\tilde{F}^{-1}_{1}(\hat{\gamma})\in\mathbb{R}^{r\times r}$, $\tilde{F}^{-1}_{2}(\hat{\gamma})\in\mathbb{R}^{(p-r)\times(p-r)}$, and $\tilde{F}^{-1}_{3}(\hat{\gamma})\in\mathbb{R}^{(K-r)\times(K-r)}$. In addition, we use $\kappa(\hat{\gamma})\in\mathbb{R}^{n\times n}$ to denote the diagonal matrix with entries $(h^{\prime}(\tilde{g}^{\top}_{i}\hat{\gamma}))^{2}/v(\tilde{g}^{\top}_{i}\hat{\gamma})$, $1\leq i\leq n$, on the diagonal:
$\kappa(\hat{\gamma})=\operatorname{diag}\left(\frac{(h^{\prime}(\tilde{g}^{\top}_{i}\hat{\gamma}))^{2}}{v(\tilde{g}^{\top}_{i}\hat{\gamma})}\right).$
@@
    $n\left(\hat{\alpha}-\frac{1}{n}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)^{\mathrm{T}}\tilde{O}\left(\hat{\alpha}-\frac{1}{n}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)\ \to\ \chi_{K-r}^{2},$ (20)
    in distribution, where $\chi_{K-r}^{2}$ denotes the $\chi^{2}$ distribution with $K-r$ degrees of freedom, and $\tilde{O}=n^{-1}\left(\kappa(\hat{\gamma})-\kappa(\hat{\gamma})\tilde{Z}\left(\tilde{Z}^{\top}\kappa(\hat{\gamma})\tilde{Z}\right)^{-1}\tilde{Z}^{\top}\kappa(\hat{\gamma})\right)$.
@@
To understand condition (21), note that according to (10), $u^{\top}\beta^{*}$ lies in the linear space spanned by coordinates of $Z_{(r+1):p}^{\top}XGu$. Condition (21) essentially requires that this space is not vanishing or degenerate. Otherwise, the inference of $u^{\top}\beta^{*}$ would not be meaningful. Condition (23) has a similar interpretation. These conditions are also needed in Le and Li (2022). Note also that in Theorem 3, $n^{-1}\|\tilde{Z}_{(r+1):p}^{T}XGu\|$, $n^{-1}\|\tilde{Z}_{1:r}^{T}XGu\|$, and $\tilde{O}$ are invariant to the choices of bases for $S_{K}(\hat{P})$ and $\operatorname{col}(X)$.
Corollary 2 and Theorem 3 provide the asymptotic distributions for $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\theta}$ that can be used for inference purposes. In particular, (20), (24), and (22) allow us to test the presence of pure network effect (against $\alpha^{*}=0$), pure covariate effect (against $\theta^{*}=0$), and the shared information between the two (against $\beta^{*}=0$), respectively. For example, when testing against $H_{0}:\alpha^{*}=0$, Theorem 3 indicates that we can use $n\hat{\alpha}^{T}\tilde{O}\hat{\alpha}$ as the statistic for a $\chi^{2}$ test.
@@ 4 Simulation Studies
We now introduce a sequence of simulation experiments to evaluate the proposed method. In particular, we want to demonstrate the effectiveness of the proposed method in model estimation and statistical inference under two types of perturbation mechanisms for network structures:random network model perturbations and embedding perturbations from deep learning embedding algorithms. We will evaluate two specific models in our model class: the subspace logistic regression and the subspace Poisson regression.
@@ 4.1 Perturbations from random network models
We first study the performance of the proposed methods when the observed networks are subject to the perturbations introduced by random network models. In particular, the true relational matrix $P$ in our model is assumed to be a probability matrix taking values in $[0,1]$. Our true model is defined based on $S_{K}(P)$. The observed network is generated from $P$ following the so-called “inhomogeneous Erdös-Rényi” framework:for each $i<j,i,j\in[n]$, generate the adjacency matrix $A_{ij}\sim\text{Bernoulli}(P_{ij})$. Different matrices $P$ tend to generate networks with different structures and the perturbation comes from the randomness of this generating process.
@@
Random network models. Regarding the network generative mechanisms, we use two low-rank models and a full-rank model. The stochastic block model (SBM) of Holland et al. (1983) with three communities, and the out-in-ratio between blocks is set to be $0.3$. The degree-corrected block model (DCBM) of Karrer and Newman (2011), where the community connection matrix is the same as the SBM with additional degree parameters varying from 0.2 to 1 (before rescaling). These two models are generated by the R package randnet (Li et al., 2023). The full-rank model is the one from Zhang et al. (2017), in which $P$ is constructed from the graphon function $g(\mu,\nu)=c/\{1+\exp[15(0.8|\mu-\nu|)^{4/5}-0.1]\}$. This grahpon model gives a banded matrix along the diagonal. So we call it the “diagonal graphon” model. In all experiments, we vary the sample size $n$ from $500$ to $4000$, and the expected average degree is set to be $\varphi_{n}=2\log n,\sqrt{n},n^{2/3},n/6$ to demonstrate the effect of varying network density.
Subsapce and covariates. Following Le and Li (2022), we construct $X\in\mathbb{R}^{n\times p}$ using the eigenvectors $w_{1},\ldots,w_{n}$ from $P$ as follows:Set $X_{1}/\sqrt{n}=w_{1};$ Set $X_{2}/\sqrt{n}=w_{2}/5+2\sqrt{6}w_{4}/5$. This configuration yields a design with $r=1,s=3$. We set $\beta=(0,0.5)^{\top}$ and $\theta=(0.5,0)^{\top}$ in all settings. Similarly, $\gamma^{*}$ can be any vector with the first coordinate being zero, where we set $\gamma^{*}=(0,0.5,0.5)^{\top}$. In the model fitting process, we always use the observed adjacency matrix $A$ to approximate the true eigenspace.
@@
Evaluation criterion. For model estimation accuracy, we measure the performance by the mean squared error (MSE) on $\beta_{2}$, defined as $|\hat{\beta}_{2}-\beta_{2}|^{2}$, the mean square prediction error (MSPE) defined as $\|\hat{Y}-\mathbb{E}Y\|^{2}/n$. For inference, we evaluate the coverage probability of the 95% confidence interval for $\beta_{2}$  [^1] The theoretical findings in Theorem 3 suggests that a valid inference cannot be drawn for $\hat{\beta}_{1}$ for the overlapping effect, similar to Le and Li (2022).
Calculation procedure. In order to assess the model’s performance with the randomness from both the response $Y$ and neighborhood matrix $A$, we generate $100$ unique neighborhood matrices $A$ based on one relational matrix $P$ for each simulation scenario. For each given $A$, $1000$ replicates of $Y$’s are generated, and the performance metrics, the coverage probability, and MSE are calculated based on the Monte Carlo approximation from these 1000 instantiations. At the outer loop, we repeatedly generate $A$ 100 times, and the median value of the resulting coverage probabilities and MSEs are reported.
@@
Table 1: Median MSE and coverage probability for subspace logistic regression under random network perturbations.
@@
Table 1 shows how our method performs under the network subspace logistic regression model. Overall, the performance improves with the same size $n$ and the expected average degree of the network model. The denser networks make the problem easier because the concentration of the adjacency matrix to the true $P$ is better. The study by Le and Li (2022) suggests that the small perturbation projection assumption holds true under random network perturbation for networks with an average degree higher than $\sqrt{n}$. Table 1 confirms this observation.
Table 2: Median MSPE (×$\times 10^{2}$) for subspace logistic regression and benchmarks under traditional random network perturbations.
@@
Table 2 presents the MSE comparison between our model and the two benchmarks. The standard logistic regression has the worst performance as it lacks network information. The RNC method captures general network information with its cohesion penalty, and its effectiveness is not very sensitive to network sparsity. Especially under the simplest SBM, it is empirically known that the penalty can accurately capture community information, as demonstrated by Li et al. (2019). Therefore, RNC can be accurate under the SBM, with advantages in sparse settings. However, when network degrees are heterogeneous (under the DCBM), RNC loses its effectiveness under the subspace model. Our method remains effective in all cases.
Under the Poisson model, we use the same configuration except for replacing the logistic distribution with the Poisson distribution. The same results are presented in Table 3 and Table 4. In Table 3, the phase transition phenomenon at the expected average degree of $\sqrt{n}$ is clearer than in Table 1. When the average degree surpasses the order of $\sqrt{n}$, the asymptotic validity holds. Under the diagonal graphon model, the perturbation has a stronger impact, but the inference remains approximately correct with the current sample size for sufficiently dense networks. The overall message remains the same as in the logistic regression setting.
Table 3: Median MSE and coverage probability for subspace Poisson regression under random network perturbations.
@@
In conclusion, the simulation results indicate that the estimator from our model is asymptotically unbiased and normally distributed when the average degree is $n^{2/3}$ or higher. However, the asymptotic behavior of estimation on sparser networks is influenced by the choice of link function and data scale.
@@
4.2 Network perturbations by from deep-learning-based embedding methods

### v6

@@ 3 Statistical Inference Properties
This section provides theoretical results for estimation consistency and statistical inference of the proposed method. To this end, we need the following regularity conditions.
@@
Assumption 3 (Boundedness of design vectors).
@@
There exist constants $\delta,C>0$ such that when $\|\gamma-\gamma^{*}\|<\delta$ the oracle information matrix defined in (13) satisfies that
@@
$\tau_{n}:=n^{-3/2}\frac{\|(\tilde{W}\tilde{W}^{\top}-WW^{\top})Z\|}{\min\left\{\left(1-\sigma_{r+1}\right)^{3},\sigma_{r+s}^{3}\right\}},$
@@
$\tau_{n}=o(n^{-1/2}).$
Assumption 5, which is also used in Le and Li (2022), is our essential requirement for the level of tolerable network perturbation. This assumption is not directly verifiable unless one specifies both the network’s perturbation mechanism (typically unknown in practice) and the choice of $\hat{P}$. For example, under the “inhomogeneous Erdős-Rényi” model, choosing the adjacency matrix $A$ as $\hat{P}$ may require a dense network (average degree above $\sqrt{n}$) for Assumption 5 to hold, as suggested by Le and Li (2022). However, Le and Li (2022) also shows that using parametric estimation to denoise $A$ can yield a $\hat{P}$ that requires a much weaker assumption under specific models. Generally speaking, one should leverage more efficient estimators of the probability matrix $P$ where appropriate to make Assumption 5 easier to hold. Notable examples include the nonparametric estimators proposed by Zhang et al. (2017) and Li and Le (2023) for general network models, as well as model-specific estimators developed in Ma et al. (2020) and Rubin-Delanchy et al. (2022). A rigorous theoretical analysis of these estimators falls outside the scope of the present work. However, readers should note that we do not restrict ourselves to the inhomogeneous Erdős-Rényi framework. Assumption 5 should be interpreted more broadly—as a robustness criterion applicable beyond a specific network generative model. In our simulation study (Section 4), for instance, we consider a scenario in which the perturbation is from a deep-learning-based embedding (which is clearly not an inhomogeneous Erdős-Rényi model) and the corresponding $\hat{P}$ is the similarity matrix of the embeddings. Empirically, we show that our method yields valid inference in this setting as well.
Assumption 6 (Moment constraints for responses).
There exist constants $c>0$, $M_{0}>0$ and $\xi>2$ such that
@@
Assumption 6 provides a sufficient condition for the Lindeberg-Feller Central Limit Theorem to hold. A similar constraint has been adopted in Yin et al. (2006); Gao et al. (2012).
@@
Consider the estimating equation (16) and assume that Assumptions 16 hold. There exists $\hat{\gamma}$ such that as $n\to\infty$,
@@
Theorem 1 shows that for each $n$, there exists a solution to the estimating equation (16) with high probability. In addition, the sequences of corresponding estimates for the true parameters in (3) are consistent. It is worth noting that similar to Yin et al. (2006), Theorem 1 itself does not guarantee the uniqueness of the solution $\hat{\gamma}$. This is because the log-likelihood function is generally not concave for certain link functions. However, Corollary 2 below shows that restricting the model space to the class with natural link functions, or more generally, link functions ensuring concavity, leads to the uniqueness.
@@
Suppose Assumptions 1 to 6 hold and the link function is natural. That is, $h^{-1}=\psi$. Then the estimates in Theorem 1 are unique for sufficiently large $n$.
[…]
@@
    $n\left(\hat{\alpha}-\frac{1}{n}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)^{\top}\tilde{O}\left(\hat{\alpha}-\frac{1}{n}\tilde{W}_{(r+1):K}\tilde{W}^{\top}_{(r+1):K}\alpha^{*}\right)\ \to\ \chi_{K-r}^{2},$ (20)
[…]
@@
[…]
Corollary 2 and Theorem 3 provide the asymptotic distributions for $\hat{\alpha}$, $\hat{\beta}$, and $\hat{\theta}$ that can be used for inference purposes. In particular, (20), (24), and (22) allow us to test the presence of pure network effect (against $\alpha^{*}=0$), pure covariate effect (against $\beta^{*}=0$), and the shared information between the two (against $\theta^{*}=0$), respectively. For example, when testing against $H_{0}:\alpha^{*}=0$, Theorem 3 indicates that we can use $n\hat{\alpha}^{\top}\tilde{O}\hat{\alpha}$ as the statistic for a $\chi^{2}$ test with KrK-r degrees of freedom.
@@ 4 Simulation Studies
We next present simulation experiments evaluating estimation and inference under two perturbation mechanisms—random‑network perturbations and embedding‑induced perturbations. We study two instances of our model: subspace logistic and subspace Poisson regression.
@@ 4.1 Perturbations from random network models
We first study the performance of the proposed methods when the observed networks are subject to the perturbations introduced by random network models. In particular, the true relational matrix $P$ in our model is assumed to be a probability matrix taking values in $[0,1]$. Our true model is defined based on $S_{K}(P)$. The observed network is generated from $P$ following the “inhomogeneous Erdös-Rényi” framework: for each $i<j,i,j\in[n]$, generate edges $A_{ij}\sim\text{Bernoulli}(P_{ij})$. Different matrices $P$ tend to generate networks with different structures and the perturbation comes from the randomness of this generating process.
@@
Random network models. Regarding the network generative mechanisms, we use two low-rank models and a full-rank model. The first is the stochastic block model (SBM) of Holland et al. (1983) with three communities, and the out-in-ratio between blocks is set to be $0.3$. The second model is the degree-corrected block model (DCBM) of Karrer and Newman (2011), where the community connection matrix is the same as the SBM with additional degree parameters varying from 0.2 to 1 (before rescaling). These two models are generated by the R package randnet (Li et al., 2023). The full-rank model is the one from Zhang et al. (2017), in which $P$ is constructed from the graphon function $g(\mu,\nu)=c/\{1+\exp[15(0.8|\mu-\nu|)^{4/5}-0.1]\}$. This graphon model gives a banded matrix along the diagonal. We therefore refer to it as the “diagonal graphon” model. In all experiments, we vary the sample size $n$ from $500$ to $4000$, and the expected average degree is set to be $\varphi_{n}=2\log n,\sqrt{n},n^{2/3}$ to demonstrate the effect of varying network density.
Subspace and covariates. Following Le and Li (2022), we construct $X\in\mathbb{R}^{n\times p}$ using the eigenvectors $w_{1},\ldots,w_{n}$ from $P$ as follows: Set $X_{1}/\sqrt{n}=w_{1};$ Set $X_{2}/\sqrt{n}=w_{2}/5+2\sqrt{6}w_{4}/5$  [^1] The main purpose of this design is to find an eigenvector that is orthogonal to $w_{1},...,w_{3}$, so we can easily control the values such as $r,s,\sigma_{r+1}$, etc. It does not have to be $w_{4}$. . This configuration yields a design with $r=1,s=1$, and the singular values of $Z^{\top}W$ in (6) are well separated: $\sigma_{1}=1,\quad\sigma_{2}=1/5,\quad\sigma_{3}=0,$ ensuring a clear distinction between signal and noise components. This separation guarantees that all regularity conditions are satisfied, except for Assumption 5, which specifically concerns network perturbation magnitude. By keeping $\min\left\{\left(1-\sigma_{r+1}\right)^{3},\sigma_{r+s}^{3}\right\}$ fixed, we can then systematically control and vary the magnitude of perturbations by the average degree of the network. We set $\beta^{*}=(0,0.5)^{\top}$ and $\theta^{*}=(0.5,0)^{\top}$ in all settings. Similarly, we set $\gamma_{3:4}^{*}=(0.5,0.5)^{\top}$. Then we generate $Y$ from the logistic regression model or Poisson regression model separately, following
$Y\mid X$ $\sim$ $\text{Bernoulli}\left\{\frac{\exp\left(X\beta^{*}+X\theta^{*}+\alpha^{*}\right)}{1+\exp\left(X\beta^{*}+X\theta^{*}+\alpha^{*}\right)}\right\},$
$Y\mid X$ $\sim$ $\text{Poisson}\left\{\exp\left(X\beta^{*}+X\theta^{*}+\alpha^{*}\right)\right\}.$
In the model fitting process, we always use the observed adjacency matrix $A$ to approximate the true eigenspace.
@@
Evaluation criterion. For model estimation accuracy, we measure the performance by the mean squared error (MSE) on $\beta_{2}$, defined as $|\hat{\beta}_{2}-\beta_{2}^{*}|^{2}$, the mean square prediction error (MSPE) defined as $\|\hat{Y}-\mathbb{E}Y\|^{2}/n$. For inference, we evaluate the coverage probability of the 95% confidence interval for $\beta_{2}$  [^2] It can be shown that, for $\beta_{1}$, where $u=(1,0)^{\top}$, we have $\left\|Z_{2}^{\top}XGu\right\|=0.$ This configuration violates the requirement $\frac{1}{n}\left\|Z_{r+1:p}^{\top}XGu\right\|\geq c$ in (21) for Theorem 3. This implies that the parameter subspace relevant for inference is degenerate. We construct this setting intentionally to examine the theoretical assumption.
@@
Table 1: Median MSE ($\times 10^{2}$) and coverage probability for subspace logistic regression under random network perturbations.
@@
Table 2: Median MSPE ($\times 10^{2}$) for subspace logistic regression and benchmarks under traditional random network perturbations.
@@
Table 3: Median MSE ($\times 10^{-2}$) for $\hat{\beta}_{2}$ and rejection rate of $\chi^{2}$ test for subspace logistic regression, under random network perturbations when $\left\|\alpha^{*}\right\|=0$.
n avg. degree SBM DCBM Diag
MSE Rejection MSE Rejection MSE Rejection
500 $2\log n$ 1.06 4.9% 1.10 4.6% 1.04 5.0%
$\sqrt{n}$ 1.05 4.8% 1.11 4.8% 1.03 4.8%
$n^{2/3}$ 1.07 4.8% 1.10 5.0% 1.00 4.8%
1000 $2\log n$ 0.51 4.8% 0.53 4.8% 0.52 4.7%
$\sqrt{n}$ 0.51 4.9% 0.53 4.8% 0.51 4.7%
$n^{2/3}$ 0.52 4.9% 0.53 5.0% 0.50 4.7%
2000 $2\log n$ 0.27 4.9% 0.27 4.9% 0.25 4.9%
$\sqrt{n}$ 0.27 4.9% 0.27 4.8% 0.24 5.0%
$n^{2/3}$ 0.27 5.0% 0.27 5.0% 0.24 5.1%
4000 $2\log n$ 0.13 4.9% 0.13 4.9% 0.12 5.0%
$\sqrt{n}$ 0.13 5.0% 0.13 5.1% 0.12 5.1%
$n^{2/3}$ 0.13 5.0% 0.13 4.8% 0.12 4.8%
Calculation procedure. In order to assess the model’s performance with the randomness from both the response $Y$ and adjacency matrix $A$, we generate $100$ unique adjacency matrices $A$ based on one relational matrix $P$ for each simulation scenario. For each given $A$, $1000$ replicates of $Y$’s are generated, and the performance metrics (MSE and MSPE) and coverage probability are computed based on the Monte Carlo approximation from these 1000 instantiations. In the outer loop, we repeatedly generate $A$ 100 times, and the median value of the resulting coverage probabilities and MSEs are reported.
Table 4: Median MSE ($\times 10^{2}$) and coverage probability for subspace Poisson regression under random network perturbations.
@@
Table 6: Median MSE ($\times 10^{-2}$) for $\hat{\beta}_{2}$ and rejection rate of $\chi^{2}$ test for subspace Poisson regression, under random network perturbations when $\left\|\alpha^{*}\right\|=0$.
n avg. degree SBM DCBM Diag
MSE Rejection MSE Rejection MSE Rejection
500 $2\log n$ 0.11 4.8% 0.11 5.0% 0.14 4.8%
$\sqrt{n}$ 0.11 5.0% 0.11 4.8% 0.14 5.2%
$n^{2/3}$ 0.11 5.0% 0.11 5.2% 0.12 5.1%
1000 $2\log n$ 0.06 4.9% 0.06 5.2% 0.08 4.9%
$\sqrt{n}$ 0.06 4.9% 0.05 5.0% 0.07 4.8%
$n^{2/3}$ 0.06 4.8% 0.05 5.1% 0.06 4.9%
2000 $2\log n$ 0.03 5.0% 0.03 5.1% 0.04 5.0%
$\sqrt{n}$ 0.03 4.8% 0.03 4.9% 0.03 4.8%
$n^{2/3}$ 0.03 4.9% 0.03 4.8% 0.03 5.1%
4000 $2\log n$ 0.01 4.9% 0.01 5.1% 0.02 5.0%
$\sqrt{n}$ 0.01 4.8% 0.01 4.9% 0.02 5.0%
$n^{2/3}$ 0.01 4.8% 0.01 5.0% 0.02 5.0%
Table 1 shows how our method performs under the network subspace logistic regression model. Overall, the performance improves with the sample size $n$ and the expected average degree of the network model. The denser networks make the problem easier because the concentration of the adjacency matrix to the true $P$ is better. Table 1 shows that if $A$ is used under the current network generative procedure, an average degree higher than $\sqrt{n}$ is sufficient for good inference accuracy. This is consistent with the observation in Le and Li (2022). Table 2 presents the MSE comparison between our model and the two benchmarks. Our method clearly outperforms the RNC and standard logistic regression.
Table 3 summarizes the model’s performance while $\left\|\alpha^{*}\right\|=0$. The specific focus is the rejection rate of the $\chi^{2}$ test at the level 0.05. The results suggest that the $\chi^{2}$ test performs well under the null hypothesis with the desired level of type I error control.
Under the Poisson model, we use the same configuration except for replacing the logistic distribution with the Poisson distribution. The same results are presented in Table 4, Table 5 and Table 6. When the average degree surpasses the order of $\sqrt{n}$, the asymptotic validity holds. Under the diagonal graphon model, the perturbation has a stronger impact, but the inference remains approximately correct with the current sample size for sufficiently dense networks. The overall message remains the same as in the logistic regression setting.
@@
4.2 Network perturbations from deep-learning-based embedding methods